# Triangular solutions to a Riemann-Hilbert problem from superelliptic curves

### v1

@@ 4.  Monodromy matrices of solutions
$\mathcal{M}_{i}=\mathrm{diag}\left(e^{2\pi\mathrm{i}\beta_{1}^{(i)}},\ldots,e^{2\pi\mathrm{i}\beta_{p}^{(i)}}\right)=\mathrm{diag}\left(e^{2\pi\mathrm{i}\theta_{i}},e^{2\pi\mathrm{i}\theta_{i}-\frac{n}{m}},\ldots,e^{2\pi\mathrm{i}\theta_{i}-\frac{(p-1)n}{m}}\right).$
@@
$\mathcal{M}_{i}={\rm e}^{2\pi\bf{\rm i}\theta_{i}}\begin{pmatrix}1&\hat{\alpha}_{1}\left(1-{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}\right)&\hat{\alpha}_{2}\left(1-{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\right)-\hat{\alpha}_{1}^{2}\left({\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}-{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\right)\\ 0&{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}&\hat{\alpha}_{1}\left({\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}-{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\right)\\ 0&0&{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\end{pmatrix}.$
@@
$\mathcal{M}_{i}={\rm e}^{2\pi\bf{\rm i}\theta_{i}}\begin{pmatrix}1&\alpha\left(1-{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}\right)&\frac{1}{2}\alpha^{2}\left(1+{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}-2{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}\right)\\ 0&{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}&\alpha\left({\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}-{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\right)\\ 0&0&{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\end{pmatrix}.$
We see that in all cases the monodromy matrices are independent of small variations of $a_{i}$, $i=1,\dots,N$, and are conjugate to ${\rm exp}\{2\pi\bf{\rm i}\,B^{(i)}\}$ as expected according to the general theory of Fuchsian systems, see for example, [16]. Note that our solutions also transform when $z$ goes around the point at infinity, and the corresponding monodromy matrix $\mathcal{M}_{\infty}$ is conjugate to ${\rm exp}\{2\pi\bf{\rm i}\,B^{(\infty)}\}$ where $B^{(\infty)}=\sum_{i=1}^{N}B^{(i)}$ is a nonzero diagonal matrix, see the end of Section 2. A loop around $z=\infty$ is given by a product of loops $\rho_{i}$ and thus $\mathcal{M}_{\infty}$ is obtained as the corresponding product of the matrices $\mathcal{M}_{i}$.
5.  Some rational solutions
As mentioned after Theorem 1, the contours of integration may be chosen in a way for the solutions of the Schlesinger system from Theorem 1 to be given by residues of the differentials (6). From [7], we know that for $n>0$, the differentials only have poles at the $s$ points at infinity $\infty_{1},\dots,\infty_{s}$. For coprime $m$ and $N$, there is only one pole at infinity and thus the residue there is zero. For $n<0$, all poles are at the finite ramification points $P_{a_{1}},\dots,P_{a_{N}}.$ In these cases we obtain rational solutions to the Schlesinger system and to the Fuchsian system as well.
@@ 5.1  Positive values of $n$
(29) $\Omega_{i}^{(j)}(P)=\frac{w^{jn}d\zeta}{\zeta-a_{i}}=\frac{\nu_{jn\alpha}\prod_{i=1}^{N}(1-a_{i}\xi^{m_{1}})^{jn/m}}{\xi^{jnN_{1}+1}(1-a_{i}\xi^{m_{1}})}\,d\xi,\qquad\nu_{jn\alpha}=-m_{1}\,e^{2\pi\bf{\rm i}jn(\alpha-1)/s}.$
@@
Let $\gamma_{\infty_{\alpha}}$ be a small contour encircling $\infty_{\alpha}$. Choosing now $\gamma_{l-k}=\frac{c_{l-k}}{2\pi\bf{\rm i}A_{l-k}}\gamma_{\infty_{\alpha}}$ for some $c_{l-k}\in\mathbb{C}$, we have a polynomial solution to the Schlesinger system given, for some $\theta_{i}\in\mathbb{C}$, by (see Theorem 3 of [7])
@@
(32) $\underset{\infty_{\alpha}}{\rm res}\,\Omega_{i}^{(l-k)}(a)=A_{l-k}\sum_{k_{1}+\ldots+k_{N}+q=N(l-k)\frac{n}{m}}(-1)^{q}{(l-k)\frac{n}{m}\choose k_{1}}\ldots{(l-k)\frac{n}{m}\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}a_{i}^{q},$
@@
(${\beta\choose j}=\frac{\beta(\beta-1)\cdots(\beta-j+1)}{j!},\qquad{\beta\choose 0}=1$
defined for $\beta\in\mathbb{R}$ and $j\in\mathbb{N}$. Solutions obtained in Theorem 4 to the Fuchsian system (1) with polynomial coefficients (31) are rational functions of $z$ given by the next corollary.
@@
Let $\theta_{i}\in\mathbb{C}$ for $i=1,\dots,N$ and $p,n,m,N$ be strictly positive integers such that ${\rm gcd}(n,m)=1$ and $s={\rm gcd}(m,N)>1$. With $c_{j}\in\mathbb{C}$ as in (31), let $R_{j}=R_{j}(z)$ be the rational funtion
(33) $R_{j}(z)=\begin{dcases*}-\frac{m\,c_{j}}{jn}\sum_{k_{1}+\ldots+k_{N}+q=N\frac{jn}{m}}(-1)^{q}{{jn}/{m}\choose k_{1}}\ldots{{jn}/{m}\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}z^{q},\quad\text{if }j/m_{1}\in\mathbb{Z}\\ 0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\;\text{if }j/m_{1}\notin\mathbb{Z}.\end{dcases*}$
Denote $\beta_{k}^{(i)}=\theta_{i}-\frac{(k-1)n}{m}$ for $k=1,\dots,p$ and $i=1,\dots,N$. Then the following upper triangular matrix $\Phi(z)$ solves the Fuchsian system (1) with coefficients $B^{(i)}$ being the upper triangular $p\times p$ matrices given by (31).
@@
(36) $M_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{R_{j}^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\quad&\text{when } $l>k$,\\ \quad 1\quad&\text{when } $l=k$,\\ \quad 0\quad&\text{when } $l<k$,\end{dcases*}$
@@
Proof.   Let $z\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$. For $n>0$ and the compact Riemann surface $X_{a}$ corresponding to the superelliptic curve (4), the differentials $\Omega_{i}^{(j)}$ (6) have poles at $\infty_{1},\dots,\infty_{s}$ as explained at the beginning of Section 5.1. Comparing the statements of Corollary 8 and Theorem 4, we only need to prove that $\kappa_{j}$ from Definition 2 coincides with $R_{j}$ (33) for the choice of the integration contour $\gamma_{j}=\frac{c_{j}}{2\pi\bf{\rm i}A_{j}}\gamma_{\infty_{\alpha}}$ with $\gamma_{\infty_{\alpha}}$ being a small contour encircling $\infty_{\alpha}\in X_{a}$ and not containing in its interior any point $P\in X_{a}$ for which $\zeta(P)=z$. With this choice and $m_{1},N_{1}$ as above, we have
@@
$\kappa_{j}$ $=$ $-\frac{m\,c_{j}}{jnA_{j}}\,\underset{\xi=0}{\rm res}\,\frac{\nu_{jn\alpha}d\xi}{\xi^{jnN_{1}+1}}\sum_{k_{1}=0}^{\infty}{jn/m\choose k_{1}}(-a_{1}\xi^{m_{1}})^{k_{1}}\ldots\sum_{k_{N}=0}^{\infty}{jn/m\choose k_{N}}(-a_{N}t^{m_{1}})^{k_{N}}\sum_{q=0}^{\infty}(z\xi^{m_{1}})^{q}$
$=$ $-\frac{m\,c_{j}}{jnA_{j}}\,\underset{\xi=0}{\rm res}\,\frac{\nu_{jn\alpha}d\xi}{\xi^{jnN_{1}+1}}\sum_{r=0}^{\infty}\Bigl{[}\sum_{k_{1}+\ldots+k_{N}+q=r}(-1)^{r-q}{jn/m\choose k_{1}}\ldots{jn/m\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}z^{q}\Bigr{]}\,\xi^{rm_{1}}.$
@@
Let us set, for simplicity, $\theta_{1}=\theta_{2}=\frac{1}{4}$, $p=2$, and $n=1,\;m=2,\;N=2$ so that $s=2$. In this case, $\beta_{1}=\frac{1}{4}$ and $\beta_{2}=-\frac{1}{4}$ and the polynomial in $a_{1},\,a_{2}$ solution (31) to the Schlesinger system is
@@ 5.2  Negative values of $n$
Thus, $\Omega_{i}^{(j)}$ has a pole of order $j|n|-m+1$ at $P_{a_{\nu}}$ with $\nu\neq i$ if $j|n|\geqslant m$ and of order $j|n|+1$ at $P_{a_{i}}.$ Note that, given that $n$ is coprime to $m$, the residues are nonzero only if $j$ is a multiple of $m$. Let us choose contours of integration in (8) in the form
(38) $\gamma_{l-k}=\sum_{\nu=1}^{N}\frac{c^{\nu}_{l-k}}{2\pi\bf{\rm i}}\gamma_{a_{\nu}}$
@@
(40) $\underset{P_{a_{\nu}}}{\rm res}\,\Omega_{i}^{(l-k)}(a)=m\sum_{k_{1}+\ldots+k_{N}=(l-k)|n|/m-1}\frac{(-1)^{k_{\nu}}}{(a_{\nu}-a_{i})^{k_{\nu}+1}}\prod_{h=1,h\neq\nu}^{N}\frac{{-(l-k)|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+(l-k)|n|/m}}$
@@
(41) $\underset{P_{a_{\nu}}}{\rm res}\,\Omega_{\nu}^{(l-k)}(a)=m\sum_{k_{1}+\ldots+\hat{k}_{\nu}+\ldots+k_{N}=(l-k)|n|/m}\prod_{h=1,h\neq\nu}^{N}\frac{{-(l-k)|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+(l-k)|n|/m}}$
where the hat symbol over $k_{\nu}$ indicates that the summation index $k_{\nu}$ is absent in the sum in (41). Solutions obtained in Theorem 4 to the Fuchsian system (1) with polynomial coefficients (39) are rational functions of $z$ given by the next corollary.
@@
Let $n<0,$ and $p,m,N>$ be integers such that ${\rm gcd}(n,m)=1$. Let $\theta_{i}\in\mathbb{C}$ for $i=1,\dots,N$ and denote $\beta_{k}^{(i)}=\theta_{i}-\frac{(k-1)n}{m}$ for $k=1,\dots,p$. With $c_{j}\in\mathbb{C}$ as in (39), let $S_{j}=S_{j}(z)$ be the rational funtion
(42) $S_{j}(z)=\begin{dcases*}\frac{-m^{2}}{jn}\sum_{\nu=1}^{N}c^{\nu}_{j}\sum_{k_{1}+\ldots+k_{N}=j|n|/m-1}\frac{(-1)^{k_{\nu}}}{(a_{\nu}-z)^{k_{\nu}+1}}\prod_{h=1,h\neq\nu}^{N}\frac{{-j|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+j|n|/m}},\quad\text{if }j/m\in\mathbb{Z}\\ \hskip 8.5359pt0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\text{if }j/m\notin\mathbb{Z}.\end{dcases*}$
@@
(45) $M_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{S_{j}^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\quad&\text{when } $l>k$,\\ \quad 1\quad&\text{when } $l=k$,\\ \quad 0\quad&\text{when } $l<k$,\end{dcases*}$
@@
Proof.   Let $z\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$. For $n<0$ and the compact Riemann surface $X_{a}$ corresponding to the superelliptic curve (4), the differentials $\Omega_{i}^{(j)}$ (6) have poles at $P_{a_{\nu}}$ for $\nu=1,\dots,N$ as shown at the beginning of Section 5.2. Similarly to the proof of Corollary 8, we need to prove that $\kappa_{j}$ from Definition 2 for the choice of the integration contour $\gamma_{j}$ as in (38) with $l-k=j$ coincides with $S_{j}$ (42). We make an additional assumption that none of the contours $\gamma_{a_{\nu}}$ from (38) contains in its interior any point $P\in X_{a}$ for which $\zeta(P)=z$. With this choice, in terms of the standard local parameter $\xi_{a_{\nu}}$ (2) at $P_{a_{\nu}}$, we have
@@
Now, to obtain the residue, for each $\nu$, we take the term of the series in $\xi_{a_{\nu}}$ with $r$ such that $rm=j|n|-m$ if such an integer $r$ exists. If $j|n|-m$ is not an integer multiple of $m$, which is equivalent to $j$ not being an integer multiple of $m$ due to the choice of $n$ coprime with $m$, the residue vanishes. This gives us $S_{j}$ from the statement of the corollary. $\Box$
@@
Let $p=3,\;m=N=2$, and $n=-1$. We thus consider the family of curves given by equation $w^{2}=(\zeta-a_{1})(\zeta-a_{2})$ where the associated Riemann surfaces are of genus zero. We define both integration contours to be a constant multiple of the small contour $\gamma_{a_{2}}$ around the point $P_{a_{2}}=(\zeta=a_{2},w=0)$ oriented counterclockwise on the surface: $\gamma_{1}=\gamma_{2}=\frac{c}{2\pi\bf{\rm i}}\gamma_{a_{2}}.$ Note that for every given $z\in\mathbb{C}$ the contour $\gamma_{a_{2}}$ is assumed not to encircle the points $P$ of the surface for which $\zeta(P)=z.$ For any $\theta_{1},\theta_{2},c\in\mathbb{C}$, Corollary 10 gives that the matrix
@@
Acknowledgments. BP would like to thank Maxime Fortier Bourque (Université de Montréal) for insightful discussions regarding sums over partitions of integers. BP also acknowledges that this research was partially funded by the Fonds de recherche du Québec - section Nature et Technologies (FRQ). VS gratefully acknowledges support from the Natural Sciences and Engineering Research Council of Canada through a Discovery grant and from the University of Sherbrooke.

### v2

@@ 4. Monodromy matrices of solutions
$\mathcal{M}_{i}=\mathrm{diag}\left(e^{2\pi{\mathrm{i}}\beta_{1}^{(i)}},\ldots,e^{2\pi{\mathrm{i}}\beta_{p}^{(i)}}\right)=\mathrm{diag}\left(e^{2\pi{\mathrm{i}}\theta_{i}},e^{2\pi{\mathrm{i}}(\theta_{i}-\frac{n}{m})},\ldots,e^{2\pi{\mathrm{i}}(\theta_{i}-\frac{(p-1)n}{m})}\right).$
@@
$\mathcal{M}_{i}={\rm e}^{2\pi{\mathrm{i}}\theta_{i}}\begin{pmatrix}1&\hat{\alpha}_{1}\left(1-{\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}\right)&\hat{\alpha}_{2}\left(1-{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}\right)-\hat{\alpha}_{1}^{2}\left({\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}-{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}\right)\\ 0&{\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}&\hat{\alpha}_{1}\left({\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}-{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}\right)\\ 0&0&{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}\end{pmatrix}$.
@@
$\mathcal{M}_{i}={\rm e}^{2\pi{\mathrm{i}}\theta_{i}}\begin{pmatrix}1&\alpha\left(1-{\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}\right)&\frac{1}{2}\alpha^{2}\left(1+{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}-2{\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}\right)\\ 0&{\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}&\alpha\left({\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}-{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}\right)\\ 0&0&{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}\end{pmatrix}$.
We see that in all cases the monodromy matrices are independent of small variations of $a_{i}$, $i=1,\dots,N$, and are conjugate to ${\rm exp}\{2\pi{\mathrm{i}}\,B^{(i)}\}$ as expected according to the general theory of non-resonant Fuchsian systems, see Chapter IV, Theorem 4.1 of [6]. Note that our solutions also transform when $z$ goes around the point at infinity, and the corresponding monodromy matrix $\mathcal{M}_{\infty}$ is conjugate to ${\rm exp}\{2\pi{\mathrm{i}}\,B^{(\infty)}\}$ where $B^{(\infty)}=\sum_{i=1}^{N}B^{(i)}$ is a nonzero diagonal matrix, see the end of Section 2. A loop around $z=\infty$ is given by a product of loops $\rho_{i}$ and thus $\mathcal{M}_{\infty}$ is obtained as the corresponding product of the matrices $\mathcal{M}_{i}$.
5. Some solutions $\Phi=MD$ with rational $M$
As mentioned after Theorem 1, the contours of integration may be chosen in a way for the solutions of the Schlesinger system from Theorem 1 to be given by residues of the differentials (6). From [8], we know that for $n>0$, the differentials only have poles at the $s$ points at infinity $\infty_{1},\dots,\infty_{s}$. For coprime $m$ and $N$, there is only one pole at infinity and thus the residue there is zero. For $n<0$, all poles are at the finite ramification points $P_{a_{1}},\dots,P_{a_{N}}.$ In these cases we obtain, respectively, polynomial and rational solutions to the Schlesinger system. For the corresponding solutions $\Phi=MD$ to the Fuchsian system, the matrix $M$ is rational. The solution $\Phi$ is algebraic for rational constants $\theta_{i}$, $i=1,\dots,N,$ from Theorem 4.
@@ 5.1  Positive values of $n$
(29) $\Omega_{i}^{(j)}(P)=\frac{w^{jn}d\zeta}{\zeta-a_{i}}=\frac{\nu_{jn\alpha}\prod_{i=1}^{N}(1-a_{i}\xi^{m_{1}})^{jn/m}}{\xi^{jnN_{1}+1}(1-a_{i}\xi^{m_{1}})}\,d\xi,\qquad\nu_{jn\alpha}=-m_{1}\,e^{2\pi{\mathrm{i}}jn(\alpha-1)/s}.$
@@
Let $\gamma_{\infty_{\alpha}}$ be a small contour encircling $\infty_{\alpha}$. Choosing now $\gamma_{l-k}=\frac{c_{l-k}}{2\pi{\mathrm{i}}A_{l-k}}\gamma_{\infty_{\alpha}}$ for some $c_{l-k}\in\mathbb{C}$, we have a polynomial solution to the Schlesinger system given, for some $\theta_{i}\in\mathbb{C}$, by (see Theorem 3 of [8])
@@
(32) $\underset{\infty_{\alpha}}{\rm res}\,\Omega_{i}^{(l-k)}(a)=A_{l-k}\sum_{k_{1}+\ldots+k_{N}+t=N(l-k)\frac{n}{m}}(-1)^{t}{(l-k)\frac{n}{m}\choose k_{1}}\ldots{(l-k)\frac{n}{m}\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}a_{i}^{t},$
@@
${\beta\choose j}=\frac{\beta(\beta-1)\cdots(\beta-j+1)}{j!},\qquad{\beta\choose 0}=1$
defined for $\beta\in\mathbb{R}$ and $j\in\mathbb{N}$. Solutions obtained in Theorem 4 to the Fuchsian system (1) with polynomial coefficients (31) are functions of $z$ given by the next corollary. These solutions are algebraic for rational values of $\theta_{i}$.
@@
Let $\theta_{i}\in\mathbb{C}$ for $i=1,\dots,N$ and $p,n,m,N$ be strictly positive integers such that ${\rm gcd}(n,m)=1$ and $s={\rm gcd}(m,N)>1$. With $c_{j}\in\mathbb{C}$ as in (31), let $R_{j}=R_{j}(z)$ be the polynomial
(33) $R_{j}(z)=\begin{dcases*}-\frac{m\,c_{j}}{jn}\sum_{k_{1}+\ldots+k_{N}+t=N\frac{jn}{m}}(-1)^{t}{{jn}/{m}\choose k_{1}}\ldots{{jn}/{m}\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}z^{t},\quad\text{if }j/m_{1}\in\mathbb{Z}\\ 0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\;\text{if }j/m_{1}\notin\mathbb{Z}.\end{dcases*}$
Denote $\beta_{k}^{(i)}=\theta_{i}-\frac{(k-1)n}{m}$ for $k=1,\dots,p$ and $i=1,\dots,N$. Then the following upper triangular matrix $\Phi(z)$ solves the Fuchsian system (1) with coefficients $B^{(i)}$ being the upper triangular $p\times p$ matrices given by (31):
@@
[…]
@@
Proof.   Let $z\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$. For $n>0$ and the compact Riemann surface $X_{a}$ corresponding to the superelliptic curve (4), the differentials $\Omega_{i}^{(j)}$ (6) have poles at $\infty_{1},\dots,\infty_{s}$ as explained at the beginning of Section 5.1. Comparing the statements of Corollary 8 and Theorem 4, we only need to prove that $\kappa_{j}$ from Definition 2 coincides with $R_{j}$ (33) for the choice of the integration contour $\gamma_{j}=\frac{c_{j}}{2\pi{\mathrm{i}}A_{j}}\gamma_{\infty_{\alpha}}$ with $\gamma_{\infty_{\alpha}}$ being a small contour encircling $\infty_{\alpha}\in X_{a}$ and not containing in its interior any point $P\in X_{a}$ for which $\zeta(P)=z$. With this choice and $m_{1},N_{1}$ as above, we have
@@
$\kappa_{j}$ $=$ $-\frac{m\,c_{j}}{jnA_{j}}\,\underset{\xi=0}{\rm res}\,\frac{\nu_{jn\alpha}d\xi}{\xi^{jnN_{1}+1}}\sum_{k_{1}=0}^{\infty}{jn/m\choose k_{1}}(-a_{1}\xi^{m_{1}})^{k_{1}}\ldots\sum_{k_{N}=0}^{\infty}{jn/m\choose k_{N}}(-a_{N}\xi^{m_{1}})^{k_{N}}\sum_{t=0}^{\infty}(z\xi^{m_{1}})^{t}$
$=$ $-\frac{m\,c_{j}}{jnA_{j}}\,\underset{\xi=0}{\rm res}\,\frac{\nu_{jn\alpha}d\xi}{\xi^{jnN_{1}+1}}\sum_{r=0}^{\infty}\Bigl{[}\sum_{k_{1}+\ldots+k_{N}+t=r}(-1)^{r-t}{jn/m\choose k_{1}}\ldots{jn/m\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}z^{t}\Bigr{]}\,\xi^{rm_{1}}.$
@@
Let us set, for simplicity, $\theta_{1}=\theta_{2}=\frac{1}{4}$, $p=2$, and n=1,m=2,N=2n=1,\;m=2,\;N=2 so that $s=2$. In this case, $\beta_{1}=\frac{1}{4}$ and $\beta_{2}=-\frac{1}{4}$ and the polynomial in $a_{1},\,a_{2}$ solution (31) to the Schlesinger system is
@@ 5.2  Negative values of $n$
Thus, $\Omega_{i}^{(j)}$ has a pole of order j|n|m+1j|n|-m+1 at $P_{a_{\nu}}$ with $\nu\neq i$ if $j|n|\geqslant m$ and of order $j|n|+1$ at $P_{a_{i}}.$ Note that, given that $n$ is coprime to $m$, the residues are nonzero only if $j$ is a multiple of $m$. Let us choose contours of integration in (8) in the form
(38) $\gamma_{l-k}=\sum_{\nu=1}^{N}\frac{c^{\nu}_{l-k}}{2\pi{\mathrm{i}}}\gamma_{a_{\nu}}$
@@
(40) $\underset{P_{a_{\nu}}}{\rm res}\,\Omega_{i}^{(l-k)}(a)=m\hskip-28.45274pt\sum_{k_{1}+\ldots+k_{N}=(l-k)|n|/m-1}\frac{(-1)^{k_{\nu}}}{(a_{\nu}-a_{i})^{k_{\nu}+1}}\prod_{h=1,h\neq\nu}^{N}\frac{{-(l-k)|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+(l-k)|n|/m}}$
@@
(41) $\underset{P_{a_{\nu}}}{\rm res}\,\Omega_{\nu}^{(l-k)}(a)=m\hskip-28.45274pt\sum_{k_{1}+\ldots+\hat{k}_{\nu}+\ldots+k_{N}=(l-k)|n|/m}\prod_{h=1,h\neq\nu}^{N}\frac{{-(l-k)|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+(l-k)|n|/m}}$
where the hat symbol over $k_{\nu}$ indicates that the summation index $k_{\nu}$ is absent in the sum in (41). Solutions obtained in Theorem 4 to the Fuchsian system (1) with rational coefficients (39) are given by the next corollary.
@@
Let $n<0,$ and $p,m,N>0$ be integers such that ${\rm gcd}(n,m)=1$. Let $\theta_{i}\in\mathbb{C}$ for $i=1,\dots,N$ and denote $\beta_{k}^{(i)}=\theta_{i}-\frac{(k-1)n}{m}$ for $k=1,\dots,p$. With $c_{j}^{\nu}\in\mathbb{C}$ as in (39), let $S_{j}=S_{j}(z)$ be the rational function
(42) $S_{j}(z)=\begin{dcases*}\frac{-m^{2}}{jn}\sum_{\nu=1}^{N}c^{\nu}_{j}\hskip-8.5359pt\sum_{k_{1}+\ldots+k_{N}=j|n|/m-1}\frac{(-1)^{k_{\nu}}}{(a_{\nu}-z)^{k_{\nu}+1}}\prod_{h=1,h\neq\nu}^{N}\frac{{-j|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+j|n|/m}},\quad\text{if }j/m\in\mathbb{Z}\\ \hskip 8.5359pt0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\text{if }j/m\notin\mathbb{Z}.\end{dcases*}$
@@
[…]
@@
Proof.   Let $z\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$. For $n<0$ and the compact Riemann surface $X_{a}$ corresponding to the superelliptic curve (4), the differentials $\Omega_{i}^{(j)}$ (6) have poles at $P_{a_{\nu}}$ for $\nu=1,\dots,N$ as shown at the beginning of Section 5.2. Similarly to the proof of Corollary 8, we need to prove that $\kappa_{j}$ from Definition 2 for the choice of the integration contour $\gamma_{j}$ as in (38) with lk=jl-k=j coincides with $S_{j}$ (42). We make an additional assumption that none of the contours $\gamma_{a_{\nu}}$ from (38) contains in its interior any point $P\in X_{a}$ for which $\zeta(P)=z$. With this choice, in terms of the standard local parameter $\xi_{a_{\nu}}$ (2) at $P_{a_{\nu}}$, we have
@@
Now, to obtain the residue, for each $\nu$, we take the term of the series in $\xi_{a_{\nu}}$ with $r$ such that rm=j|n|mrm=j|n|-m if such an integer $r$ exists. If j|n|mj|n|-m is not an integer multiple of $m$, which is equivalent to $j$ not being an integer multiple of $m$ due to the choice of $n$ coprime with $m$, the residue vanishes. This gives us $S_{j}$ from the statement of the corollary. $\Box$
@@
Let p=3,m=N=2p=3,\;m=N=2, and n=1n=-1. We thus consider the family of curves given by equation $w^{2}=(\zeta-a_{1})(\zeta-a_{2})$ where the associated Riemann surfaces are of genus zero. We define both integration contours to be a constant multiple of the small contour $\gamma_{a_{2}}$ around the point $P_{a_{2}}=(\zeta=a_{2},w=0)$ oriented counterclockwise on the surface: $\gamma_{1}=\gamma_{2}=\frac{c}{2\pi{\mathrm{i}}}\gamma_{a_{2}}.$ Note that for every given $z\in\mathbb{C}$ the contour $\gamma_{a_{2}}$ is assumed not to encircle the points $P$ of the surface for which $\zeta(P)=z.$ For any $\theta_{1},\theta_{2},c\in\mathbb{C}$, Corollary 10 gives that the matrix
@@
Acknowledgments. BP would like to thank Maxime Fortier Bourque (Université de Montréal) for insightful discussions regarding sums over partitions of integers. BP also acknowledges that this research was partially funded by the Fonds de recherche du Québec - section Nature et Technologies (FRQ). The authors thank Renat Gontsov for careful reading of the manuscript and pointing out some errors of notation and terminology, which greatly improved the presentation. VS gratefully acknowledges support from the Natural Sciences and Engineering Research Council of Canada through a Discovery grant and from the University of Sherbrooke.
Data availability statement. The authors declare that there is no data associated with the manuscript.
Conflict of interest statement. On behalf of all authors, the corresponding author states that there is no conflict of interest.